# Stacked Intelligent Metasurface for End-to-End OFDM System

### v1

@@ Abstract
Stacked intelligent metasurface (SIM) and dual-polarized SIM (DPSIM) enabled wave-domain signal processing have emerged as promising research directions for offloading baseband digital processing tasks and efficiently simplifying transceiver design. However, existing architectures are limited to employing SIM (DPSIM) for a single communication function, such as precoding or combining. To further enhance the overall performance of SIM (DPSIM)-assisted systems and achieve end-to-end (E2E) joint optimization from the transmitted bitstream to the received bitstream, we propose an SIM (DPSIM)- assisted E2E orthogonal frequency-division multiplexing (OFDM) system, where traditional communication tasks such as modulation, precoding, combining, and demodulation are performed simultaneously during electromagnetic (EM) forward propagation. Furthermore, inspired by the idea of abstracting real metasurfaces as hidden layers of a neural network, we propose the electromagnetic neural network (EMNN) to enable the control of the E2E OFDM communication system. In addition, transfer learning is introduced into the model training, and a training and deployment framework for the EMNN is designed. Simulation results demonstrate that both SIM-assisted E2E OFDM systems and DPSIM-assisted E2E OFDM systems can achieve robust bitstream transmission under complex channel conditions. Our study highlights the application potential of EMNN and SIM (DPSIM)-assisted E2E OFDM systems in the design of next-generation transceivers.
Index Terms:
Stacked intelligent metasurfaces (SIM), Dual-polarized stacked intelligent metasurfaces (DPSIM), Holographic multiple-input multiple-output (HMIMO), Orthogonal frequency-division multiplexing (OFDM), electromagnetic neural network (EMNN), Reconfigurable intelligent metasurface (RIS), End-to-end (E2E).
I Introduction
@@
Recently, reconfigurable intelligent surfaces (RISs) have emerged as a pivotal technology capable of directly manipulating electromagnetic (EM) waves in the wave domain, thereby potentially offloading a substantial portion of the signal processing tasks traditionally handled by RF chains [8]. Building upon this concept, a novel antenna architecture has been proposed in which the RIS is directly integrated into the transceiver [9, 10]. By precisely controlling the EM responses of the RIS elements, the transceiver can perform joint signal processing in both the digital and wave domains. This integrated architecture, referred to as holographic multiple-input multiple-output (HMIMO) [11], effectively reduces the number of required RF chains and typically employs low-resolution digital-to-analog converters (DACs) and analog-to-digital converters (ADCs), thereby significantly lowering hardware costs and improving energy efficiency [12].
However, due to the inherent single-layer structure of RISs, their flexibility in manipulating EM waves remains limited [13]. To further enhance wave-domain signal processing capabilities, [14, 15] proposed HMIMO systems assisted by stacked intelligent metasurfaces (SIMs), which significantly outperform conventional massive MIMO systems and RIS-assisted HMIMO systems. Specifically, a SIM is a three-dimensional metasurface device that leverages cascaded multilayer metasurfaces to perform analog computing in the EM wave domain [16]. In practice, SIMs are typically integrated into the RF front end to adjust the amplitude and phase of the transmitted or received signals, thereby replacing the precoding (or combining) modules in traditional architectures and ultimately simplifying transceiver design [17]. Furthermore, [18] proposed stacking dual-polarized RISs to enable independent signal processing in two orthogonal polarization directions, thereby further improving the signal processing capability within the constraints of limited integration space.
@@
SIM and DPSIM: Since the works in [14, 15] first proposed utilizing SIMs to accomplish signal processing tasks in the wave domain, a series of follow-up studies on SIMs have emerged. The authors of [25] investigated the application of SIMs in multi-user beamforming and achieved higher spectral efficiency and lower energy consumption compared with conventional MIMO systems. The study in [26] explored the potential of SIMs in wideband HMIMO systems, enabling interference-free multicarrier transmission under frequency-selective fading. In [27], a hybrid optimization method was proposed to achieve rate maximization and minimize inter-symbol interference (ISI), demonstrating that increasing the number of metasurface layers can improve the overall system performance. The work in [28] developed a multiport network model for SIMs based on Z-parameters to address issues such as unidirectional propagation and neglected mutual coupling in conventional simplified approaches. Furthermore, SIMs have been applied to various domains, including integrated sensing and communication (ISAC) [29] and semantic communications [30]. Research on dual-polarized SIMs (DPSIMs) is also thriving.  [18] first proposed a DPSIM-assisted HMIMO system and demonstrated that, under the same integration space, the DPSIM-assisted system outperforms the SIM-assisted system in terms of performance. In [31], a DPSIM transceiver architecture combined with rate-splitting (RS-DPSIM) was proposed, where common and private information are transmitted simultaneously along two polarization directions. Simulation results show that this architecture offers performance advantages in anti-interference communications.
@@
E2E learning of OFDM system: In recent years, extensive research has focused on exploring the performance of E2E learning in various communication systems. In [32], an autoencoder (AE) was introduced into orthogonal frequency-division multiplexing (OFDM) systems for the first time, assuming that each subcarrier is processed independently, thus enabling E2E communication for OFDM. However, this approach restricts the receiver’s ability to exploit the time–frequency correlation of the OFDM channel. To address this limitation, [33] designed and optimized a neural network receiver based on fully connected layers to jointly process multiple subcarriers, significantly improving system performance. The work in [34] investigated the impact of quantization bit resolution on E2E communication in OFDM systems and verified the robustness of learning-based methods in this scenario. Moreover, [35] studied E2E learning with OFDM waveforms over frequency- and time-selective fading channels, demonstrating the performance gains achieved through joint pilot and constellation learning. In addition, novel network architectures, such as the generative adversarial network (GAN) [23] and the deep fully convolutional neural network (DeepRx) [36], have also been applied to E2E communications to further enhance system performance.
@@
    We established the mathematical models of the SIM-assisted E2E OFDM systems and the DPSIM-assisted E2E OFDM systems. Different from existing studies on SIM (DPSIM), the proposed system aims to achieve joint optimization of the communication link from the transmitted bit stream to the received bit stream. By deploying deep neural networks (DNNs) and integrating SIM (DPSIM) devices, the systems enable complete offloading of signal processing tasks from the baseband digital domains to the wave domains.
@@
    To achieve holistic optimization of the E2E system performance, we further propose the electromagnetic neural network (EMNN) by abstracting the physical metasurface as hidden layers in the network. The EMNN associates the hidden-layer parameters with the EM (DPEM) units, thereby enabling model-driven control.
@@
    The effectiveness of employing transfer learning to accelerate the training of EMNN is validated through simulations. In addition, a comprehensive performance evaluation of the SIM (DPSIM)-assisted E2E OFDM system is conducted, and its performance is compared with that of conventional massive MIMO systems.
@@
II System Model
Figure 2: TX-SIM/RX-SIM device parameter diagram.
@@ II System Model
As shown in Fig. 1, we consider a $J$-user E2E OFDM system, where transmit SIM (TX-SIM) is integrated in the base station (BS) transmitter, and receive SIM (RX-SIM) is integrated in each user equipment (UE) receiver. To simplify the model, we assume that all UEs have the same number of receiving antennas and each integrates an identical RX-SIM device. Furthermore, all metasurface layers constituting a SIM device are identical. Let $B$ and $N^{\mathrm{c}}$ denote the bandwidth and the number of subcarriers of the OFDM system, respectively. Moreover, as shown in Fig. 2, $A^{\mathrm{t}}=A^{\mathrm{t}^{\mathrm{x}}}\times A^{\mathrm{t}^{\mathrm{y}}}$ and $A^{\mathrm{r}}=A^{\mathrm{r}^{\mathrm{x}}}\times A^{\mathrm{r}^{\mathrm{y}}}$ respectively denote the number of TX antennas and the number of RX antennas per user. We further assume that TX-SIM comprises $L$ layers of metasurfaces, with $M=M^{\mathrm{x}}\times M^{\mathrm{y}}$ EM units on each layer, and one RX-SIM is equipped with $K$ layers of metasurfaces, with $N=N^{\mathrm{x}}\times N^{\mathrm{y}}$ EM units on each layer. For convenience of notation, we define the set of TX metasurfaces as $\mathcal{L}=\left\{0,1,2,\cdots,L\right\}$ and the set of RX metasurfaces for the $j$-th user as $\mathcal{K}_{j}=\left\{0,1,2,\cdots,K\right\}$, where $0\in\mathcal{L}$ and $0\in\mathcal{K}_{j}$ respectively represent the BS’s TX antenna metasurface and the $j$-th user’s RX antenna metasurface. The spacing between adjacent EM units in the TX-SIM and RX-SIM is denoted as $d^{\mathrm{t}}$ and $d^{\mathrm{r}}$, respectively.
Figure 3: Schematic diagram of single-polarization multipath channel model.
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \boldsymbol{\alpha}_{i,s}^{\mathrm{t}}=\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{x}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{t}}\sin\mathrm{(}\vartheta_{s}^{\mathrm{t}})\sin\mathrm{(}\varphi_{s}^{\mathrm{t}})\right)\otimes\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{y}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{t}}\cos\mathrm{(}\vartheta_{s}^{\mathrm{t}})\right),$ (10)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \boldsymbol{\alpha}_{i,j,s}^{\mathrm{r}}=\boldsymbol{\alpha}_{j,s}^{\mathrm{r}^{\mathrm{x}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{r}}\sin\mathrm{(}\vartheta_{j,s}^{\mathrm{r}})\sin\mathrm{(}\varphi_{j,s}^{\mathrm{r}})\right)\otimes\boldsymbol{\alpha}_{j,s}^{\mathrm{r}^{\mathrm{y}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{r}}\cos\mathrm{(}\vartheta_{j,s}^{\mathrm{r}})\right),$ (11)
@@
$\ \ \ \ \ \ \mathbf{\Phi}_{p}^{l}=\mathrm{diag}\left(e^{\mathrm{j}\theta_{p,1}^{l}},e^{\mathrm{j}\theta_{p,2}^{l}},\cdots,e^{\mathrm{j}\theta_{p,M}^{l}}\right)\in\mathbb{C}^{M^{\mathrm{dp}}\times M^{\mathrm{dp}}},\theta_{p,m}^{l}\in\left[0,2\pi\right),m\in\mathcal{M}^{\mathrm{dp}},l\in\mathcal{L}^{\mathrm{dp}},p\in\left\{0,1\right\},$ (16)
$\ \ \ \ \ \ \mathbf{\Psi}_{j,p}^{k}=\mathrm{diag}\left(e^{\mathrm{j}\xi_{p,1}^{j,k}},e^{\mathrm{j}\xi_{p,2}^{j,k}},\cdots,e^{\mathrm{j}\xi_{p,N}^{j,k}}\right)\in\mathbb{C}^{N^{\mathrm{dp}}\times N^{\mathrm{dp}}},\xi_{p,n}^{j,k}\in\left[0,2\pi\right),n\in\mathcal{N}^{\mathrm{dp}},k\in\mathcal{K}^{\mathrm{dp}},p\in\left\{0,1\right\},1\leqslant j\leqslant J.$ (17)
@@
where $\mathbf{n}_{i,j}\in\mathbb{C}^{A^{\mathrm{r}}\times 1}$ is the receiver noise vector with distribution $\mathcal{C}\mathcal{N}\left(0,\sigma^{2}\mathbf{I}_{A^{\mathrm{r}}}\right)$, $\mathbf{x}_{i}\in\mathbb{C}^{A^{\mathrm{t}}\times 1}$ is the signal vector satisfying $\mathbb{E}\left\{\mathbf{x}_{i}\mathbf{x}_{i}^{\mathrm{H}}\right\}=\mathbf{I}_{A^{\mathrm{t}}}$, $\mathbf{p}_{i}\in\mathbb{C}^{A^{\mathrm{t}}\times 1}$ denotes the vector of transmit powers, $\mathbf{G}_{i,j}$ is the channel coefficient matrix between the $L$-th layer metasurface of the TX-SIM and the $K$-th layer metasurface of the $j$-th user’s RX-SIM.
Let $N^{\mathrm{bit}}$ denote the number of bits corresponding to an OFDM symbol transmitted by the BS, and $N_{j}^{\mathrm{bit}}$ represent the number of bits received by the $j$-th user within one OFDM symbol. $\mathbf{b}\in\{0,1\}^{N^{\mathrm{bit}}\times 1}$ and $\hat{\mathbf{b}}_{j}\in\{0,1\}^{N_{j}^{\mathrm{bit}}\times 1}$ denote the transmitted bit vector and the received bit vector, respectively. Then, the SIM-assisted E2E OFDM system model can be expressed as
@@
II-B Unpolarized Channel Model (SIM-assisted E2E OFDM System Channel Model) .
As shown in Fig. 3, to precisely model the spatial and frequency-selective properties inherent in wideband propagation, the proposed SIM-assisted E2E OFDM system utilizes a geometric multi-path channel model. Assume there are $S$ scatterers between the BS and UEs, each of which affects the signal propagation path. Let $s=0$ denote the line-of-sight (LoS) path and $s=1,2,\cdots,S$ denote non-LoS (NLoS) paths. We assume that the gain and delay of the $i$-th subcarrier from the BS to the $j$-th user through the $s$-th path are $g_{i,j,s}$ and $\tau_{i,j,s}$, respectively. Although the transmission signals between TX and RX are time-domain signals, the wideband channel model is often represented in the frequency-domain form to simplify the IFFT-FFT progress in the beamspace[38]. For the $i$-th subcarrier with frequency $f_{i}$, the channel coefficient matrix $\mathbf{G}_{i,j}\in\mathbb{C}^{N\times M}$ is modeled as
@@
Without loss of generality, we assume that the TX-SIM and RX-SIM are vertically aligned in a 3D Cartesian coordinate system. For the $s$-th propagation path, its elevation and azimuth angles relative to the TX-SIM are $\vartheta_{s}^{\mathrm{t}}\in[0,\pi/2)$ and $\varphi_{s}^{\mathrm{t}}\in[0,2\pi)$, respectively, while those relative to the $j$-th user are $\vartheta_{j,s}^{\mathrm{r}}\in[0,\pi/2)$ and $\varphi_{j,s}^{\mathrm{r}}\in[0,2\pi)$.
The steering vector $\boldsymbol{\alpha}_{i,s}^{\mathrm{t}}$ and $\boldsymbol{\alpha}_{i,j,s}^{\mathrm{r}}$ can be written as (10) and (11), respectively, where $\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{x}}}(*)\in\mathbb{C}^{M^{\mathrm{x}}\times 1}$ and $\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{y}}}(*)\in\mathbb{C}^{M^{\mathrm{y}}\times 1}$ are defined as follows

### v2

@@ Abstract
Stacked intelligent metasurface (SIM) and dual-polarized SIM (DPSIM) enabled wave-domain signal processing have emerged as promising research directions for offloading baseband digital processing tasks and efficiently simplifying transceiver design. However, existing architectures are limited to employing SIM (DPSIM) for a single communication function, such as precoding or combining. To further enhance the overall performance of SIM (DPSIM)-assisted systems and achieve end-to-end (E2E) joint optimization from the transmitted bitstream to the received bitstream, we propose an SIM (DPSIM)-assisted E2E orthogonal frequency division multiplexing (OFDM) system, in which traditional communication tasks such as channel coding, modulation, precoding, combining, demodulation, and channel decoding are performed synchronously within the electromagnetic (EM) forward propagation. Furthermore, inspired by the idea of abstracting real metasurfaces as hidden layers of a neural network, we propose the EM neural network (EMNN) to enable the control of the E2E OFDM communication system. In addition, transfer learning is introduced into the model training, and a training and deployment framework for the EMNN is designed. Simulation results demonstrate that both SIM-assisted E2E OFDM systems and DPSIM-assisted E2E OFDM systems can achieve robust bitstream transmission under complex channel conditions. Our study highlights the application potential of EMNN and SIM (DPSIM)-assisted E2E OFDM systems in the design of next-generation transceivers.
I Introduction.
@@
Recently, reconfigurable intelligent surfaces (RISs) have emerged as a pivotal technology capable of directly manipulating electromagnetic (EM) waves in the wave domain, thereby potentially offloading a substantial portion of the signal processing tasks traditionally handled by RF chains [8]. Building upon this concept, a novel antenna architecture has been proposed in which the RIS is directly integrated into the transceiver [9, 10]. By precisely controlling the EM responses of the RIS elements, the transceiver can perform joint signal processing in both the digital and wave domains. This integrated architecture, referred to as holographic MIMO (HMIMO) [11], effectively reduces the number of required RF chains and typically employs low-resolution digital-to-analog converters (DACs) and analog-to-digital converters (ADCs), thereby significantly lowering hardware costs and improving energy efficiency [12].
However, due to the inherent single-layer structure of RISs, their flexibility in manipulating EM waves remains limited [13]. To further enhance wave-domain signal processing capabilities, [14, 15] proposed HMIMO systems assisted by stacked intelligent metasurfaces (SIMs), which significantly outperform conventional massive MIMO systems and RIS-assisted HMIMO systems. Specifically, a SIM is a three-dimensional metasurface device that leverages cascaded multilayer metasurfaces to perform analog computing in the EM wave domain [16]. In practice, SIMs are typically integrated into the RF front end to adjust the amplitude and phase of the transmitted or received signals, thereby replacing the precoding (or combining) modules in traditional architectures and ultimately simplifying transceiver design [17]. Furthermore,  [18] proposed stacking dual-polarized RISs to enable independent signal processing along two orthogonal polarization directions. This architecture, referred to as the dual-polarized SIM (DPSIM), further enhances signal processing capability under the constraint of limited integration space.
@@
SIM: Since the works in [14, 15] first proposed utilizing SIMs to accomplish signal processing tasks in the wave domain, a series of follow-up studies on SIMs have emerged. The authors of [25] investigated the application of SIMs in multi-user beamforming and achieved higher spectral efficiency and lower energy consumption compared with conventional MIMO systems. The study in [26] explored the potential of SIMs in wideband HMIMO systems, enabling interference-free multicarrier transmission under frequency-selective fading. In [27], a hybrid optimization method was proposed to achieve rate maximization and minimize inter-symbol interference, demonstrating that increasing the number of metasurface layers can improve the overall system performance. The work in [28] developed a multiport network model for SIMs based on Z-parameters to address issues such as unidirectional propagation and neglected mutual coupling in conventional simplified approaches. Furthermore, SIMs have been applied to various domains, including integrated sensing and communication (ISAC) [29] and semantic communications [30].
DPSIM: As an improved version of SIM, DPSIM is constructed by stacking DPRISs and can independently process signals along two orthogonal polarization directions, thereby achieving stronger signal-processing capability than SIM.  [18] first proposed a DPSIM-assisted HMIMO system and evaluated the performance gains achievable within the same integration footprint. Furthermore, [31] proposes a DPSIM-assisted transceiver architecture combined with rate-splitting (RS), which simultaneously transmits common and private information along two polarization directions. Simulations ultimately demonstrate the performance advantages of this architecture in anti-jamming communications.
@@
E2E learning of OFDM system: In recent years, extensive research has focused on exploring the performance of E2E learning in various communication systems.  [32] was the first to introduce autoencoders (AEs) into orthogonal frequency-division multiplexing (OFDM) systems. However, it assumes independent mapping on each subcarrier, which limits the receiver’s ability to exploit the time–frequency correlation of the channel. To address this issue,  [33] optimized a neural-network receiver to jointly process multiple subcarriers, significantly improving system performance.  [34] investigated the impact of quantized bits on the E2E communication performance of OFDM systems and verified the robustness of learning-based approaches under various complex channel conditions.  [35] studied E2E communication with OFDM waveforms over frequency- and time-selective fading channels, demonstrating performance gains from jointly learning pilots and constellation points. In addition, several novel network architectures—such as generative adversarial networks (GANs)  [23] and deep fully convolutional neural networks (DeepRx)  [36] have also been applied to E2E communications to simplify transceiver design.
@@
    We are the first to establish mathematical models for SIM-assisted E2E OFDM systems and DPSIM-assisted E2E OFDM systems. Different from existing studies on SIM (DPSIM), the proposed system aims to achieve joint optimization of the communication link from the transmitted bit stream to the received bit stream. By deploying deep neural networks (DNNs) and integrating SIM (DPSIM) devices, the systems enable complete offloading of signal processing tasks from the baseband digital domains to the wave domains.
@@
    To achieve holistic optimization of the E2E system performance, we further propose the EM neural network (EMNN) by abstracting the physical metasurface as hidden layers in the network. The EMNN associates the hidden-layer parameters with the EM (DPEM) units, thereby enabling model-driven control.
@@
    The simulation results validate the effectiveness of the transfer learning approach in accelerating EMNN training. Furthermore, we provide a comprehensive performance evaluation of SIM (DPSIM)- assisted E2E OFDM systems and compare their performance with that of conventional massive MIMO systems.
@@
II System Model.
Figure 2: TX-SIM and RX-SIM device parameter diagram.
@@
As shown in Fig. 1, we consider a $J$ user E2E OFDM system, where transmit SIM (TX-SIM) is integrated in the base station (BS) transmitter, and receive SIM (RX-SIM) is integrated in each user equipment (UE) receiver. To simplify the model, we assume that all UEs have the same number of receiving antennas and each integrates an identical RX-SIM device. Furthermore, all metasurface layers constituting a RX-SIM device are identical. Let $B$ and $N^{\mathrm{c}}$ denote the bandwidth and the number of subcarriers of the OFDM system, respectively. Moreover, as shown in Fig. 2, $A^{\mathrm{t}}=A^{\mathrm{t}^{\mathrm{x}}}\times A^{\mathrm{t}^{\mathrm{y}}}$ and $A^{\mathrm{r}}=A^{\mathrm{r}^{\mathrm{x}}}\times A^{\mathrm{r}^{\mathrm{y}}}$ respectively denote the number of TX antennas and the number of RX antennas per UE. We further assume that TX-SIM comprises $L$ layers of metasurfaces, with $M=M^{\mathrm{x}}\times M^{\mathrm{y}}$ EM units on each layer, and one RX-SIM is equipped with $K$ layers of metasurfaces, with $N=N^{\mathrm{x}}\times N^{\mathrm{y}}$ EM units on each layer. For convenience of notation, we define the set of TX metasurfaces as $\mathcal{L}=\left\{0,1,2,\cdots,L\right\}$ and the set of RX metasurfaces for the $j$-th user as $\mathcal{K}_{j}=\left\{0,1,2,\cdots,K\right\}$, where $0\in\mathcal{L}$ and $0\in\mathcal{K}_{j}$ respectively represent the BS’s TX antenna metasurface and the $j$-th user’s RX antenna metasurface. The spacing between adjacent EM units in the TX-SIM and RX-SIM is denoted as $d^{\mathrm{t}}$ and $d^{\mathrm{r}}$, respectively.
Figure 3: Schematic diagram of multipath channel model.
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \boldsymbol{\alpha}_{i,s}^{\mathrm{t}}=\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{x}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{t}}\sin\mathrm{(}\vartheta_{s}^{\mathrm{t}})\cos\mathrm{(}\varphi_{s}^{\mathrm{t}})\right)\otimes\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{y}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{t}}\sin\mathrm{(}\vartheta_{s}^{\mathrm{t}})\sin\mathrm{(}\varphi_{s}^{\mathrm{t}})\right),$ (10)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \boldsymbol{\alpha}_{i,j,s}^{\mathrm{r}}=\boldsymbol{\alpha}_{j,s}^{\mathrm{r}^{\mathrm{x}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{r}}\sin\mathrm{(}\vartheta_{j,s}^{\mathrm{r}})\cos\mathrm{(}\varphi_{j,s}^{\mathrm{r}})\right)\otimes\boldsymbol{\alpha}_{j,s}^{\mathrm{r}^{\mathrm{y}}}\left(\frac{2\mathrm{\pi c}}{f_{i}}d^{\mathrm{r}}\sin\mathrm{(}\vartheta_{j,s}^{\mathrm{r}})\sin\mathrm{(}\varphi_{j,s}^{\mathrm{r}})\right).$ (11)
@@
$\ \ \ \ \mathbf{\Phi}_{p}^{l}=\mathrm{diag}\left(e^{\mathrm{j}\theta_{p,1}^{l}},e^{\mathrm{j}\theta_{p,2}^{l}},\cdots,e^{\mathrm{j}\theta_{p,M^{\mathrm{dp}}}^{l}}\right)\in\mathbb{C}^{M^{\mathrm{dp}}\times M^{\mathrm{dp}}},\theta_{p,m}^{l}\in\left[0,2\pi\right),m\in\mathcal{M}^{\mathrm{dp}},l\in\mathcal{L}^{\mathrm{dp}},p\in\left\{0,1\right\},$ (16)
$\ \ \ \ \mathbf{\Psi}_{j,p}^{k}=\mathrm{diag}\left(e^{\mathrm{j}\xi_{p,1}^{j,k}},e^{\mathrm{j}\xi_{p,2}^{j,k}},\cdots,e^{\mathrm{j}\xi_{p,N^{\mathrm{dp}}}^{j,k}}\right)\in\mathbb{C}^{N^{\mathrm{dp}}\times N^{\mathrm{dp}}},\xi_{p,n}^{j,k}\in\left[0,2\pi\right),n\in\mathcal{N}^{\mathrm{dp}},k\in\mathcal{K}^{\mathrm{dp}},p\in\left\{0,1\right\},1\leqslant j\leqslant J.$ (17)
@@
where $\mathbf{n}_{i,j}\in\mathbb{C}^{A^{\mathrm{r}}\times 1}$ is the receiver noise vector with distribution $\mathcal{C}\mathcal{N}\left(0,\sigma^{2}\mathbf{I}_{A^{\mathrm{r}}}\right)$, $\mathbf{p}_{i}\in\mathbb{C}^{A^{\mathrm{t}}\times 1}$ denotes the vector of transmit powers, $\mathbf{G}_{i,j}$ is the channel coefficient matrix between the $L$-th layer metasurface of the TX-SIM and the $K$-th layer metasurface of the $j$-th user’s RX-SIM, $\mathbf{x}_{i}\in\mathbb{C}^{A^{\mathrm{t}}\times 1}$ is the signal vector satisfying $\mathbb{E}\left\{\mathbf{xx}^{\mathrm{H}}\right\}=\mathbf{I}_{A^{\mathrm{t}}N^{\mathrm{c}}}$, with $\mathbf{x}=\left[\mathbf{x}_{1}^{\mathrm{T}},\mathbf{x}_{2}^{\mathrm{T}},\cdots,\mathbf{x}_{N^{\mathrm{c}}}^{\mathrm{T}}\right]^{\mathrm{T}}$.
Let $N^{\mathrm{bit}}$ denote the number of bits corresponding to an OFDM symbol transmitted by the BS, and $N_{j}^{\mathrm{bit}}$ represent the number of bits received by the $j$-th user within one OFDM symbol. $\mathbf{b}\in\{0,1\}^{N^{\mathrm{bit}}\times 1}$ and $\hat{\mathbf{b}}_{j}\in\{0,1\}^{N_{j}^{\mathrm{bit}}\times 1}$ denote the transmitted bit vector and the $j$-th user’s received bit vector, respectively. Then, the bit transmission model of SIM-assisted E2E OFDM system model can be expressed as
@@
II-B Single-polarization channel model (SIM-assisted E2E OFDM system channel model).
As shown in Fig. 3, to precisely model the spatial and frequency-selective properties inherent in wideband propagation, the proposed SIM-assisted E2E OFDM system utilizes a geometric multi-path channel model. There are $S$ scatterers between the BS and UEs, each of which affects the signal propagation path. Let $s=0$ denote the line-of-sight (LoS) path and $s=1,2,\cdots,S$ denote non-LoS (NLoS) paths. We assume that the gain and delay of the $i$-th subcarrier from the BS to the $j$-th user through the $s$-th path are $g_{i,j,s}$ and $\tau_{i,j,s}$, respectively. Although the transmission signals between TX and RX are time-domain signals, the wideband channel model is often represented in the frequency-domain form to simplify the IFFT-FFT progress in the beamspace [38]. For the $i$-th subcarrier with frequency $f_{i}$, the channel coefficient matrix $\mathbf{G}_{i,j}\in\mathbb{C}^{N\times M}$ is modeled as
@@
Without loss of generality, we assume that the TX-SIM and RX-SIM are vertically aligned in a 3D cartesian coordinate system. For the $s$-th propagation path, its elevation and azimuth angles relative to the TX-SIM are $\vartheta_{s}^{\mathrm{t}}\in[0,\pi/2)$ and $\varphi_{s}^{\mathrm{t}}\in[0,2\pi)$, respectively, while those relative to the $j$-th user’s RX-SIM are $\vartheta_{j,s}^{\mathrm{r}}\in[0,\pi/2)$ and $\varphi_{j,s}^{\mathrm{r}}\in[0,2\pi)$. The steering vector $\boldsymbol{\alpha}_{i,s}^{\mathrm{t}}$ and $\boldsymbol{\alpha}_{i,j,s}^{\mathrm{r}}$ can be written as (10) and (11), respectively, where $\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{x}}}(*)\in\mathbb{C}^{M^{\mathrm{x}}\times 1}$ and $\boldsymbol{\alpha}_{s}^{\mathrm{t}^{\mathrm{y}}}(*)\in\mathbb{C}^{M^{\mathrm{y}}\times 1}$ are defined as follows